OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, five divided by three, negative five divided by eighteen, ... Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (anonymous):

Try to find a closed form expression for the \(n\)-th term in the sequence. Hint: it's a geometric sequence, find the common ratio.

OpenStudy (anonymous):

@SithsAndGiggles i dont understand how to even start it

OpenStudy (anonymous):

First term = 60 Divide by -6 (i.e. multiply by -1/6) and you get the second term = -10. Divide again by -6 (or multiply by -1/6) and you get the third term = (-10)/(-6) = 5/3. Thus the common ratio is \(-\dfrac{1}{6}\). There's a certain trait about common ratios that satisfy \(|r|<1\) that will help you figure out the rest.

OpenStudy (anonymous):

Diverges Converges; 11100 Converges; 72 Converges; 0 so how do i determine if one of these is the answer ? @SithsAndGiggles

OpenStudy (anonymous):

If the ratio satisfies \(|r|<1\), then the sequence converges. Does \(-\dfrac{1}{6}\) satisfy this?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

What is \(\left|-\dfrac{1}{6}\right|\)?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Right, and 1/6 is less than 1. So this sequence converges.

OpenStudy (anonymous):

converges at what ?

OpenStudy (anonymous):

Think about what's happening to the each next number in the sequence. You're making the value smaller and smaller. As an example, what happens if you were to cut a pie in half and ate one of the halves, then again cut in half, and again and again and so on. What would you be left with at the end of all this cutting?

OpenStudy (anonymous):

This scenario basically illustrates the geometric sequence \(\left\{1,\dfrac{1}{2},\dfrac{1}{4},\dfrac{1}{6},\cdots\right\}\). The terms get smaller and smaller, but never become negative. This sequence converges to 0.